OpenStudy (anonymous):

Please and thank you! The correct answer is: B Can anyone show me how to do it :) What is the equation of the circle with radius 9 and center (–2, 3)? A x2 + y2 – 4x + 6y + 4 = 0 B x2 + y2 + 4x – 6y – 68 = 0 C x2 + y2 + 6x – 4y + 4 = 0 D x2 + y2 – 6x + 4y – 68 = 0

OpenStudy (cwrw238):

use standard form (x-a)^2 + (y-b)^2 = r^2 where (a,b) is center and r = radius of the circle plug in your values and expand ing will give you B

OpenStudy (kainui):

So what you have to do here is complete the square for both the x and y parts of the equation to get it in the form: (x-a)^2+(y-b)^2=c^2

OpenStudy (anonymous):

and how do u get 68

OpenStudy (anonymous):

when you actually expand it out you will have 4+9-81(ignoring all the variables)

OpenStudy (anonymous):

i never got -81

OpenStudy (cwrw238):

(x + 2)^2 + (y -3)^2 = 81 right?

OpenStudy (anonymous):

ohh whoops i forgot to square 9 thanks

OpenStudy (cwrw238):

giving 81 -4 - 9 = 68